# Exploring the Frontier of Vision-Language Models: A Survey of Current Methodologies and Future Directions

### v1

@@ 2.2 Text generation with Multimodal Input
VideoCOCA Yan et al. (2022): VideoCoCa is an adaptation of the Contrastive Captioners CoCa Yan et al. (2022) model for video-text tasks. Utilizing CoCa’s generative and contrastive attentional pooling layers, VideoCoCa achieves state-of-the-art results in zero-shot video classification and text-to-video retrieval with minimal additional training. The model processes uniformly sampled frames through CoCa’s image encoder, creating a tensor representing the entire video sequence. This tensor undergoes attention-pooling layers for both generative and contrastive modeling tasks. VideoCoCa demonstrates proficiency in various video-based tasks, including video reasoning and action recognition, but faces challenges in subtle temporal relationships. Various adaptation strategies and lightweight finetuning approaches were explored, with the attentional pooler method proving most effective. The model was tested on multiple datasets, exhibiting significant improvements over the CoCa baseline. VideoCoCa consistently outperforms CoCa across various scales and tasks, showcasing its robust performance in video-text modeling.
@@
TinyGPT-V Yuan et al. (2023a): TinyGPT-V addresses the challenges posed by closed-source and computationally demanding multimodal models like GPT-4V. This model achieves high performance with lower computational requirements, utilizing a 24G GPU for training and an 8G GPU or CPU for inference. Integrating Phi-2 and pre-trained vision modules from BLIP-2 or CLIP, TinyGPT-V excels in benchmarks such as visual question-answering and referring expression comprehension, showcasing competitive performance against larger models like LLAVA in various benchmarks. The model’s compact and efficient design, combining a small backbone with large model capabilities, marks a significant step towards practical, high-performance multimodal language models for diverse applications.
@@
ChatBridge Zhao et al. (2023b): ChatBridge is a multimodal language model aiming to create versatile AI models capable of understanding diverse real-world modalities. Using language as a bridge, the model leverages language-paired two-modality data to connect different modalities. ChatBridge extends the zero-shot capabilities of large language models, undergoing a two-stage training process aligning each modality with language and fine-tuning with a new multimodal instruction dataset (MULTIS). The model demonstrates strong results on zero-shot multimodal tasks, encompassing text, image, video, and audio. However, limitations include challenges in understanding long-range videos and audio, suggesting a need for a more precise temporal modeling approach. The framework can be extended to include additional modalities like sketch and point cloud. While frozen modules alleviate the computational burden, they may lead to insufficient performance and introduce biases from pre-trained models.
@@
Macaw LLM Lyu et al. (2023): Macaw-LLM is a novel multi-modal large language model that integrates visual, audio, and textual information. It comprises a modality module for encoding multi-modal data, a cognitive module utilizing pre-trained LLMs, and an alignment module harmonizing diverse representations. The alignment module bridges multi-modal features to textual features, simplifying adaptation. A large-scale multi-modal instruction dataset is created for multi-turn dialogue. The paper acknowledges limitations, including concerns about the evaluation accurately reflecting Macaw-LLM’s capabilities. The model is not optimized for multi-turn dialogues, and potential issues like hallucination, toxicity, and fairness are not evaluated due to the unavailability of suitable evaluation suites.
@@
GPT4Tools Yang et al. (2023a): GPT4Tools, aiming to enable open-source LLMs, such as LLaMA and OPT, to efficiently use multimodal tools. It addresses challenges posed by proprietary LLMs like ChatGPT and GPT-4, which often rely on inaccessible data and high computational costs. GPT4Tools employs self-instruction to generate an instruction-following dataset, allowing open-source LLMs to solve visual problems through Low-Rank Adaptation (LoRA) optimization. The approach significantly improves tool invocation accuracy and enables zero-shot capacity for unseen tools. However, the explicit and fixed prompt approach reduces computational efficiency, prompting the exploration of implicit tool invocation methods. Despite limitations, GPT4Tools is considered a viable approach for equipping language models with multimodal tools.
@@
PandaGPT Su et al. (2023): PandaGPT is an approach enhancing large language models with visual and auditory instruction-following capabilities. PandaGPT excels in tasks like image description, video-inspired story writing, and answering audio-related questions. It seamlessly handles multimodal inputs, connecting visual and auditory information. By combining ImageBind’s multimodal encoders and Vicuna’s large language models, PandaGPT requires only aligned image-text pairs for training and exhibits emergent cross-modal behaviors for various data modalities. The paper suggests improvements, including using additional alignment data, exploring fine-grained feature extraction, generating richer multimedia content, creating new benchmarks, and addressing common language model deficiencies. Despite these considerations, PandaGPT represents a promising step toward building Artificial General Intelligence for holistic perception across diverse modalities.
@@
Table 2: Table showing a comparative analysis of various VLMs on the MME Benchmark Fu et al. (2023).XB: The model is of X billion parameters.
@@
mPLUG-Owl Ye et al. (2023a): mPLUG-Owl is a training paradigm equipping LLMs with multi-modal abilities through modularized learning of a foundation LLM, a visual knowledge module, and a visual abstractor module. The two-stage training method aligns image and text, leveraging LLM assistance while maintaining generation abilities. Experimental results demonstrate mPLUG-Owl’s superior performance in instruction and visual understanding, multi-turn conversation, and knowledge reasoning. The model exhibits unexpected abilities like multi-image correlation and multilingual understanding but has limitations, including challenges in multi-image correlation, limited multilingual training, and mixed performance in OCR of complex scenes. The model also shows potential in vision-only document comprehension, with strengths in tasks like movie review writing and code generation but limitations in other applications, indicating further exploration opportunities in document understanding and downstream applications.
@@
Ying-VLM Li et al. (2023e): Ying-VLM is trained on M3IT dataset . Models trained with M3IT show success in following human instructions, providing engaging responses, and achieving strong generalization on unseen video and Chinese tasks. The analysis indicates that increasing task numbers improves performance, and instruction diversity influences results.M3IT comprises 2.4 million instances and 400 manually written task instructions across 40 tasks.
@@
BLIVA Hu et al. (2023): BLIVA is a novel multimodal Language Learning Model designed to handle text-rich visual questions, integrating query and patch embeddings. It outperforms existing VLMs like GPT-4 and Flamingo, showing significant improvements in OCR-VQA and Visual Spatial Reasoning benchmarks. BLIVA’s architecture includes a Q-Former for instruction-aware visual features and a fully connected projection layer for additional visual information. It demonstrates a 17.72% overall improvement in the multimodal LLM benchmark (MME) compared to baseline InstructBLIP and performs well in real-world scenarios like processing YouTube thumbnail question-answer pairs.
@@
LLAVA-phi zhu2024llava: LLaVA-Phi is a highly efficient multi-modal assistant powered by the compact language model, Phi-2. This model demonstrates significant progress in compact multi-modal systems, showing that even smaller models with 2.7B parameters can effectively engage in complex dialogues blending text and visuals, given proper training. LLaVA-Phi excels in various benchmarks covering visual comprehension, reasoning, and knowledge-based perception, suggesting its suitability for real-time interaction scenarios like embodied agents. Notably, it underscores the potential of smaller language models to achieve sophisticated levels of understanding and interaction while maintaining resource efficiency. The training process involves two stages: (1) feature alignment, where a pretrained vision encoder is connected to a language model using a subset of the LAION-CC-SBU dataset, and (2) visual instruction tuning, using a combination of GPT-generated multimodal instruction-following data and VQA data to teach the model to follow multimodal instructions.
@@
MoE-LLaVA Lin et al. (2024): MoE-LLaVA is a novel training strategy for Large Vision-Language Models developed by researchers from Peking University, Sun Yat-sen University, FarReel Ai Lab, Tencent Data Platform, and Peng Cheng Laboratory. This strategy, called MoE-tuning, efficiently handles performance degradation in multi-modal learning and model sparsity by activating only the top-k experts through routers during deployment. With 3 billion sparsely activated parameters, MoE-LLaVA achieves comparable or superior performance to state-of-the-art models while reducing hallucinations in model outputs. The architecture involves a vision encoder, visual projection layer (MLP), word embedding layer, stacked LLM blocks, and MoE blocks. MoE-tuning comprises three stages: MLP training, parameter training excluding the Vision Encoder, and initializing experts in MoE followed by training only MoE layers. Evaluation on various visual understanding datasets demonstrates MoE-LLaVA’s efficiency and effectiveness, with extensive ablation studies and visualizations illustrating its effectiveness and offering insights for future research in multi-modal learning systems.
@@
Yi-VL Chen et al. (2023c): Yi Vision Language (Yi-VL) is an open-source multimodal model based on the Yi Large Language Model series, excelling in content comprehension, recognition, and multi-round conversations about images. It leads in recent benchmarks, including English and Chinese. Key features include multi-round text-image conversations, bilingual support, strong image comprehension, and fine-grained resolution at 448 $\times$× 448. Yi-VL utilizes the LLaVA architecture, comprising a vision transformer, projection module, and large language model. However, it has limitations such as supporting only visual question answering, accepting a single image input, and potential content generation issues and object identification inaccuracies in complex scenes. Additionally, it operates at a fixed resolution of 448x448, which may result in information loss for low-resolution images and a lack of additional knowledge for higher resolutions.
@@
Moondream vikhyatk (2024): Moondream is a 1.6 billion parameter model, meticulously crafted by Vikhyatk, emerges from the fusion of SigLIP, Phi-1.5, and the expansive LLaVa training dataset. Representing a significant milestone in AI research, this model is purposefully unveiled for scholarly exploration, underlining its exclusivity for non-commercial endeavors. This amalgamation of cutting-edge techniques and robust data sets underscores a commitment to advancing the frontiers of artificial intelligence, setting a new benchmark for computational prowess and innovation in the field.
@@
Shikra Chen et al. (2023e): Shikra is a Multimodal Large Language Model designed to bridge the gap in human-like referential abilities within dialogue. Shikra can handle spatial coordinate inputs and outputs in natural language, featuring a simple architecture comprising a vision encoder, alignment layer, and LLM. It eliminates the need for extra vocabularies or external plugin models, allowing seamless integration of referential dialogue tasks with various vision-language tasks. Shikra demonstrates promising performance in tasks like REC, PointQA, Image Captioning, and VQA, enabling applications such as providing object coordinates and comparing user-pointed regions. However, it currently supports only English and lacks user-friendliness for non-English speakers. Future work aims to make Shikra multilingual and explore improved coordinate representations for dense object detection and segmentation tasks. Additionally, like most LLMs, Shikra may generate harmful or counterfactual responses.
@@
BuboGPT Zhao et al. (2023c): BuboGPT is a VLM with visual grounding capabilities, designed to enhance cross-modal interaction between vision, audio, and language. It provides a fine-grained understanding of visual objects and other modalities, enabling specific object localization in images during response generation. BuboGPT incorporates an off-the-shelf visual grounding module based on SAM for entity extraction and mask correspondence in images, alongside a two-stage training scheme and instruction dataset for comprehensive text-image-audio understanding. Despite limitations such as language hallucination and insufficient capacities in Grounding QA, BuboGPT demonstrates impressive multi-modality understanding and visual grounding abilities, promising advancements in multi-modal LLMs.
@@
ChatSpot: Zhao et al. (2023a): ChatSpot has been introduced as a unified end-to-end multimodal large language model designed to enhance human-AI interaction. It supports diverse interactive forms such as mouse clicks, drag-and-drop, and drawing boxes, providing users with a flexible and seamless interactive experience. The model is built on precise referring instructions, utilizing various reference representations like points and boxes to focus on specific regions of interest. Additionally, a multi-grained vision-language instruction-following dataset is created for training ChatSpot. Experimental results demonstrate its robustness in region referring, showing minimal instances of region referring hallucination even in the presence of box noises. This highlights ChatSpot’s capability for precise region referencing and its potential to improve interactive accuracy and efficiency in multimodal large language models.
@@
MiniGPT5 Zheng et al. (2023): MiniGPT-5 introduces an innovative interleaved vision-and-language generation technique, utilizing "generative vokens" to harmonize image-text outputs. Its distinctive two-staged training strategy focuses on description-free multimodal generation, eliminating the need for comprehensive image descriptions. MiniGPT-5 enhances model integrity with classifier-free guidance, resulting in substantial improvements over baseline models like Divter on the MMDialog dataset. It consistently delivers superior or comparable multimodal outputs in human evaluations on the VIST dataset, demonstrating its efficacy across diverse benchmarks.
DRESS Chen et al. (2023b): DRESS is a large vision language model that leverages Natural Language Feedback from Large Language Models to enhance its alignment and interactions, addressing key limitations in existing LVLMs. It introduces a novel categorization of NLF into critique and refinement types to better align with human preferences and improve interaction ability in multi-turn conversations. Critique NLF identifies strengths and weaknesses, aligning LVLMs with human preferences, while refinement NLF offers concrete improvement suggestions for refining responses. To address the non-differentiable nature of NLF, conditional reinforcement learning is generalized for training. Experimental results show that DRESS generates more helpful, honest, and harmless responses, and effectively learns from feedback during multi-turn interactions compared to state-of-the-art LVLMs.
@@
X-InstructBLIP Panagopoulou et al. (2023): X-InstructBLIP is a cross-modality framework built upon frozen large language models that integrates various modalities without extensive customization. High-quality instruction tuning data is collected automatically, enabling fine-tuning for different modalities. The model performs comparably to leading-edge counterparts without extensive pre-training or customization. A novel evaluation task, Discriminative Cross-modal Reasoning (DisCRn), has been introduced to assess the model’s cross-modal abilities across disparate input modalities. X-InstructBLIP demonstrates emergent cross-modal reasoning despite separate optimization of each modality and outperforms strong captioning baselines across all examined modalities in DisCRn. However, complexities and unanswered questions within each modality highlight challenges and opportunities for future exploration both across and within modalities.
@@
VILA Lin et al. (2023b): VILA, a Visual Language model family, emerges from an enhanced pre-training recipe that systematically augments LLMs towards VLMs. VILA consistently outperforms state-of-the-art models like LLaVA1.5 across main benchmarks, showcasing its superior performance without additional complexities. Notably, VILA’s multi-modal pre-training unveils compelling properties such as multi-image reasoning, enhanced in-context learning, and improved world knowledge, marking significant advancements in visual language modeling.
@@
Table 3: Comparative analysis of leading models on 4 zero-shot video QA datasets from Li et al. (2023c). Results are reported with two tokens for each frame.

### v2

@@ 2.2 Text generation with Multimodal Input
VideoCOCA [96]: VideoCoCa is an adaptation of the Contrastive Captioners CoCa [96] model for video-text tasks. Utilizing CoCa’s generative and contrastive attentional pooling layers, VideoCoCa achieves state-of-the-art results in zero-shot video classification and text-to-video retrieval with minimal additional training. The model processes uniformly sampled frames through CoCa’s image encoder, creating a tensor representing the entire video sequence. This tensor undergoes attention-pooling layers for both generative and contrastive modeling tasks. VideoCoCa demonstrates proficiency in various video-based tasks, including video reasoning and action recognition, but faces challenges in subtle temporal relationships. Various adaptation strategies and lightweight finetuning approaches were explored, with the attentional pooler method proving most effective. The model was tested on multiple datasets, exhibiting significant improvements over the CoCa baseline. VideoCoCa consistently outperforms CoCa across various scales and tasks, showcasing its robust performance in video-text modeling.
@@
TinyGPT-V [106]: TinyGPT-V addresses the challenges posed by closed-source and computationally demanding multimodal models like GPT-4V. A notable achievement of this model is its impressive performance while utilizing minimal computational resources, requiring only 24GB for training and 8GB for inference. TinyGPT-V, after integrating Phi-2 and vision modes from CLIP, demonstrates competitive performance across various visual question answering and comprehension benchmark datasets when compared to larger models like LLAVA.The model’s compact and efficient design, combining a small backbone with large model capabilities, marks a significant step towards practical, high-performance multimodal language models for diverse applications.
@@
ChatBridge [114]: ChatBridge is a multimodal language model aiming to create versatile AI models capable of understanding diverse real-world modalities. The model utilizes language as a conduit, harnessing language-paired data to forge connections between diverse modalities. ChatBridge expands upon the zero-shot capabilities of large language models through a two-stage training procedure, aligning each modality with language and refining with a fresh multimodal instruction dataset . The model demonstrates strong results on zero-shot multimodal tasks, encompassing text, image, video, and audio. Nevertheless, there are constraints in effectively comprehending lengthy videos and audio, indicating a requirement for a more refined temporal modeling method. There is potential to expand the framework by incorporating supplementary modalities such as sketches and point clouds. While employing frozen modules helps mitigate computational constraints, it may also result in inadequate performance and introduce biases inherited from pre-trained models.
@@
Macaw LLM [59]: Macaw-LLM represents a pioneering multi-modal large language model, seamlessly blending visual, audio, and textual data. Its architecture includes a dedicated modality module for encoding multi-modal information, a cognitive module leveraging pre-trained LLMs, and an alignment module harmonizing disparate representations. This alignment facilitates the integration of multi-modal features with textual information, streamlining adaptation processes. Additionally, a comprehensive multi-modal instruction dataset has been curated to support multi-turn dialogue. However, the paper acknowledges certain limitations, particularly regarding the accuracy of the evaluation in fully capturing the capabilities of Macaw-LLM. The model is not optimized for multi-turn dialogues, and potential issues like hallucination, toxicity, and fairness are not evaluated due to the unavailability of suitable evaluation suites.
@@
GPT4Tools [98]: GPT4Tools, aiming to enable open-source LLMs, such as LLaMA and OPT, to efficiently use multimodal tools. It addresses challenges posed by proprietary LLMs like ChatGPT and GPT-4, which often rely on inaccessible data and high computational costs. GPT4Tools creates instructional datasets that support large open-source models such as LLAMA in tackling visual challenges through LORA optimization. The approach significantly improves tool invocation accuracy and enables zero-shot capacity for unseen tools. However, the explicit and fixed prompt approach reduces computational efficiency, prompting the exploration of implicit tool invocation methods. Despite limitations, GPT4Tools is considered a viable approach for equipping language models with multimodal tools.
@@
PandaGPT [74]: PandaGPT is an approach enhancing large language models with visual and auditory instruction-following capabilities. PandaGPT excels in tasks like image description, video-inspired story writing, and answering audio-related questions. It seamlessly handles multimodal inputs, connecting visual and auditory information. By combining ImageBind’s multimodal encoders and Vicuna’s large language models, PandaGPT requires only aligned image-text pairs for training and exhibits emergent cross-modal behaviors for various data modalities. The paper suggests improvements, including using additional alignment data, exploring fine-grained feature extraction, generating richer multimedia content, creating new benchmarks, and addressing common language model deficiencies. Despite these considerations, PandaGPT represents a promising step toward building Artificial General Intelligence for holistic perception across diverse modalities.
@@
Table 2: Table showing a comparative analysis of various VLMs on the MME Benchmark [28].XB: The model is of X billion parameters.
@@
mPLUG-Owl [101]: mPLUG-Owl introduces a unique training approach that empowers LLMs with multi-modal capabilities by modularizing the learning process into three key components: a foundational LLM, a visual knowledge module, and a visual abstractor module. Through a two-stage training methodology, this paradigm effectively aligns image and text data, harnessing the support of LLMs while preserving their generation capabilities. Experimental results demonstrate mPLUG-Owl’s superior performance in instruction and visual understanding, multi-turn conversation, and knowledge reasoning. The model exhibits unexpected abilities like multi-image correlation and multilingual understanding but has limitations, including challenges in multi-image correlation, limited multilingual training, and mixed performance in OCR of complex scenes. The model also shows potential in vision-only document comprehension, with strengths in tasks like movie review writing and code generation but limitations in other applications, indicating further exploration opportunities in document understanding and downstream applications.
@@
Ying-VLM [47]: Ying-VLM is trained on M3IT dataset . Models trained with M3IT show success in following human instructions, providing engaging responses, and achieving strong performance on unseen videos and tasks in the Chinese language. The analysis indicates that increasing task numbers improves performance, and instruction diversity influences results.M3It consists of 2.4 million instances, including meticulously crafted task instructions spanning across forty different task.
@@
BLIVA [34]: BLIVA is a novel multimodal Language Learning Model designed to handle text-rich visual questions, integrating query and patch embeddings. It outperforms existing VLMs like GPT-4 and Flamingo, showing significant improvements in OCR-VQA and Visual Spatial Reasoning benchmarks. BLIVA’s architecture includes a Q-Former for instruction-aware visual features and a fully connected projection layer for additional visual information. It demonstrates an overall improvement in the multimodal LLM benchmark (MME) of 17.72% compared to InstructBLIP and performs well in real-world scenarios like processing YouTube thumbnail question-answer pairs.
@@
LLAVA-phi [121]: LLaVA-Phi is a highly efficient multi-modal assistant powered by the compact language model, Phi-2. This model demonstrates significant progress in compact multi-modal systems, showing that even smaller models with 2.7B parameters can effectively engage in complex dialogues blending text and visuals, given proper training. LLaVA-Phi excels in various benchmarks covering visual comprehension, reasoning, and knowledge-based perception, suggesting its suitability for real-time interaction scenarios like embodied agents.Significantly, it highlights how smaller language models can attain advanced levels of comprehension and engagement without sacrificing resource efficiency. The training process involves two stages: (1) feature alignment, where a pretrained vision encoder is connected to a language model using a subset of the LAION-CC-SBU dataset, and (2) visual instruction tuning, using a combination of GPT-generated multimodal instruction-following data and VQA data to teach the model to follow multimodal instructions.
@@
MoE-LLaVA [49]: MoE-LLaVA is a groundbreaking training strategy for Large Vision-Language Models. Known as MoE-tuning, this innovative approach efficiently manages performance degradation in multi-modal learning and model sparsity by activating only the top k experts during deployment via routers. Despite its architecture comprising 3 billion sparsely activated parameters, MoE-LLaVA achieves comparable or superior performance to state-of-the-art models while minimizing hallucinations in model outputs. Its architecture includes a visual encoder, visual projection layer in the form of MLP, word embedding layer, stacked LLM and MoE blocks. MoE-tuning comprises three stages: MLP training, parameter training excluding the Vision Encoder, and initializing experts in MoE followed by training only MoE layers. Evaluation on various visual understanding datasets demonstrates MoE-LLaVA’s efficiency and effectiveness, with extensive ablation studies and visualizations illustrating its effectiveness and offering insights for future research in multi-modal learning systems.
@@
Yi-VL [14]: Yi Vision Language (Yi-VL) is an open-source multimodal model based on the Yi Large Language Model series, excelling in content comprehension, recognition, and multi-round conversations about images. It leads in recent benchmarks, including English and Chinese. Key features include multi-round text-image conversations, bilingual support, strong image comprehension, and fine-grained resolution at 448 $\times$× 448. Yi-VL utilizes the LLaVA architecture, comprising a vision transformer, projection module, and large language model. However, it has limitations such as supporting only visual question answering, accepting a single image input, and potential content generation issues and object identification inaccuracies in complex scenes. Additionally, it operates at a fixed resolution of 448x448, which may result in information loss for low-resolution images and a lack of additional knowledge for higher resolutions.
@@
Moondream [84]: Moondream is a 1.6 billion parameter model, meticulously crafted by Vikhyatk, emerges from the fusion of SigLIP, Phi-1.5, and the expansive LLaVa training dataset. Representing a significant milestone in AI research, this model is purposefully unveiled for scholarly exploration, underlining its exclusivity for non-commercial endeavors. This amalgamation of cutting-edge techniques and robust data sets underscores a commitment to advancing the frontiers of artificial intelligence, setting a new benchmark for computational prowess and innovation in the field.
@@
Shikra [16]: Shikra is a Multimodal Large Language Model designed to bridge the gap in human-like referential abilities within dialogue. Shikra boasts the capability to interpret spatial coordinates through natural language, facilitated by its straightforward architecture consisting of a vision encoder, alignment layer, and LLM. Unlike other models, Shikra doesn’t require additional vocabularies or external plugins, allowing for effortless integration of referential dialogue tasks with diverse vision-language tasks. Its performance is notably strong across various tasks such as REC, PointQA, Image Captioning, and VQA, enabling functionalities like offering object coordinates and comparing user-pointed regions. However, it currently supports only English and lacks user-friendliness for non-English speakers. Future work aims to make Shikra multilingual and explore improved coordinate representations for dense object detection and segmentation tasks. Additionally, like most LLMs, Shikra may generate harmful or counterfactual responses.
@@
BuboGPT [115]: BuboGPT stands out as a Visual Language Model (VLM) equipped with visual grounding capabilities, aimed at enhancing cross-modal interaction across vision, audio, and language domains. It offers a detailed comprehension of visual elements and other modalities, enabling precise localization of objects in images during response generation. Employing an off-the-shelf visual grounding module based on SAM for entity extraction and mask correspondence in images, alongside a two-stage training approach and an extensive instruction dataset, BuboGPT strives for comprehensive understanding of text-image-audio interactions.Despite facing challenges such as language hallucination and limited capacities in Grounding QA, BuboGPT exhibits remarkable proficiency in understanding multiple modalities and visual grounding tasks, signaling promising advancements in the realm of multi-modal Language Models.
@@
ChatSpot: [113]: ChatSpot has been introduced as a unified end-to-end multimodal large language model designed to enhance human-AI interaction. It supports diverse interactive forms such as mouse clicks, drag-and-drop, and drawing boxes, providing users with a flexible and seamless interactive experience. The model is built on precise referring instructions, utilizing various reference representations like points and boxes to focus on specific regions of interest. Additionally, a multi-grained vision-language instruction-following dataset is created for training ChatSpot. Experimental results demonstrate its robustness in region referring, showing minimal instances of region referring hallucination even in the presence of box noises. This highlights ChatSpot’s capability for precise region referencing and its potential to improve interactive accuracy and efficiency in multimodal large language models.
@@
MiniGPT5 [116]: MiniGPT-5 introduces an innovative interleaved vision-and-language generation technique, utilizing ”generative vokens” to harmonize image-text outputs. Its distinctive two-staged training strategy focuses on description-free multimodal generation, eliminating the need for comprehensive image descriptions. MiniGPT-5 enhances model integrity with classifier-free guidance, resulting in substantial improvements over baseline models like Divter on the MMDialog dataset. It consistently produces superior multimodal results in human assessments conducted on the VIST dataset, showcasing its effectiveness across a range of benchmarks.
DRESS [13]:DRESS, a sophisticated multimodal language model, leverages Natural Language Feedback (NLF) from Language Models (LLMs) to enhance alignment through interactive engagements, effectively mitigating key limitations seen in current Virtual Language Models (VLMs). NLF is classified into two categories: critique and refinement, aimed at closely aligning with human preferences and bolstering the model’s proficiency in multi-turn conversations. Refinement NLF provides constructive suggestions to enhance responses, while critique NLF aids in aligning VLM outputs with human preferences. Through Reinforcement Learning, DRESS is trained to handle the non-differentiable nature of NLF. Empirical findings demonstrate that DRESS facilitates the generation of more beneficial and benign outputs and adeptly learns from feedback during multi-turn interactions, surpassing state-of-the-art VLMs.
@@
X-InstructBLIP [65]: X-InstructBLIP is a cross-modality framework built upon frozen large language models that integrates various modalities without extensive customization. High-quality instruction tuning data is collected automatically, enabling fine-tuning for different modalities. The model performs comparably to leading-edge counterparts without extensive pre-training or customization. A novel evaluation task, Discriminative Cross-modal Reasoning (DisCRn), has been introduced to assess the model’s cross-modal abilities across disparate input modalities. X-InstructBLIP demonstrates emergent cross-modal reasoning despite separate optimization of each modality and outperforms strong captioning baselines across all examined modalities in DisCRn. However, complexities and unanswered questions within each modality highlight challenges and opportunities for future exploration both across and within modalities.
@@
VILA [50]: VILA, a Visual Language model family, emerges from an enhanced pre-training recipe that systematically augments LLMs towards VLMs. VILA consistently outperforms state-of-the-art models like LLaVA1.5 across main benchmarks, showcasing its superior performance without additional complexities. Notably, VILA’s multi-modal pre-training unveils compelling properties such as multi-image reasoning, enhanced in-context learning, and improved world knowledge, marking significant advancements in visual language modeling.
@@
Table 3: Comparative analysis of leading models on 4 zero-shot video QA datasets from [44]. Results are reported with two tokens for each frame.